• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yoomin

    Log entry time 08:42:33 on February 10, 2009

    Entry number 260219

    keyword=Owl Shift Summary

    Owl Shift Summary Feb 10, 2009
    
    Workers : Yoomin Oh, Al Tobias, and Navaphon Muangma
    
    00:00 Taking elastic runs smoothly
    02:20 Beam half-wave plate out
    04:20 Hall C getting the beam, MCC tuning beam, empty target in position
    05:40 CODA died, rebooted.
    06:02 L-HRS p0 set to 1GeV
    06:15 Started to take QE data.
    06:40 Controlled access for:
               polarity change(L-HRS, - to +),
               L-HRS angle to 46.22 degree(Techs),
               BigBite sieve in(Brad), and
               Compton detector work(Alexandre)
    ------------------------------------
    Run List
    
    Pol. He3 Elastic
    Run  Time        #Events Comment
    1207 XX:XX-00:07 1M
    1208 00:08-00:23 1M
    1209 00:24-00:39 1M
    1210 00:40-00:55 1M
    1211 00:56-01:11 1M
    1212 01:12-01:27 1M
    1213 01:28-01:43 1M
    1214 01:48-02:02 1M
    1215 02:03-02:18 1M
    1216 02:25-02:39 1M      Beam 1/2 plate OUT from this run!
    1217 02:40-02:56 1M
    1218 02:57-03:11 1M
    1219 03:12-03:27 1M
    1220 03:28-03:43 1M
    1221 03:44-03:59 1M
    1223 04:45-04:56 650k    BB HV trip
    1224 05:00-05:16 1M
    1225 05:17-05:32 1M
    1227 05:47-06:02 1M
    
    Pol. He3 Quasi-elastic(LHRS P0=1.0GeV)
    1229 06:16-06:35 1M
    1230 06:36-06:47 680k    Beam off for controlled access